# ReWeaver: Towards Simulation-Ready and Topology-Accurate Garment Reconstruction

### v1

@@ Abstract
High-quality 3D garment reconstruction plays a crucial role in mitigating the sim-to-real gap in applications such as digital avatars, virtual try-on and robotic manipulation. However, existing garment reconstruction methods typically rely on unstructured representations, such as 3D Gaussian Splats, struggling to provide accurate reconstructions of garment topology and sewing structures. As a result, the reconstructed outputs are often unsuitable for high-fidelity physical simulation. We propose ReWeaver, a novel framework for topology-accurate 3D garment and sewing pattern reconstruction from sparse multi-view RGB images. Given as few as four input views, ReWeaver predicts seams and panels as well as their connectivities in both the 2D UV space and the 3D space. The predicted seams and panels align precisely with the multi-view images, yielding structured 2D–3D garment representations suitable for 3D perception, high-fidelity physical simulation, and robotic manipulation. To enable effective training, we construct a large-scale dataset GCD-TS, comprising multi-view RGB images, 3D garment geometries, textured human body meshes and annotated sewing patterns. The dataset contains over 100,000 synthetic samples covering a wide range of complex geometries and topologies. Extensive experiments show that ReWeaver consistently outperforms existing methods in terms of topology accuracy, geometry alignment and seam-panel consistency.
[Uncaptioned image]Ground TruthReWeaver
@@ 2.2 Optimization-based garment reconstruction
Some optimization-based methods, as well as hybrid approaches that combine learning-based prediction with physics-guided optimization, have made notable progress in recovering simulation-ready garments by integrating differentiable physics into the reconstruction process. For example, DiffAvatar [27] leverages the XPBD [33] algorithm to drive a differentiable physics engine that simulates cloth behavior, allowing gradient-based updates to align the simulated garment with the reconstructed scan geometry. To ensure that the optimization results are plausible both physically and design-wise, and to reduce the complexity of the optimization process, [27] introduce a control cage representation on 2D patterns. However, this method relies on a predefined garment library, which limits its ability to generalize to diverse or unseen pattern structures. Dress-1-to-3 [26] addresses a more challenging setting by leveraging generative models as priors and using the CIPC [23] simulation algorithm to reconstruct simulation-ready garments from a single input image. Its initial garment patterns are estimated from a single image using SewFormer [30], but these patterns have low geometric accuracy and may even contain incorrect topology. PhysAvatar [45] combines inverse rendering and inverse physics to automatically estimate the human shape and appearance, as well as the fabric parameters of clothing, from multi-view video. However, it only reconstructs a coarse mesh that is noisy, lacks the compactness of pattern-based representations, and cannot be directly used in industrial pipelines or for easy editing.
@@ 3 Method
We introduce ReWeaver, a framework that reconstructs accurate geometry and topology of 3D garments from multi-view images with as few as four views. In the rest of this section, we first formally define the garment reconstruction task in Section 3.1, then introduce the building blocks of ReWeaver: a multi-view visual encoder adopted from VGGT [40] (Section 3.2); a consecutive module that predicts 3D curve and patch using a bi-path transformer (Section 3.3); and an intra-surface attention module that implicitly learns a flattening process that converts the 3D curves and patches into their 2D counterparts (Section 3.4). FFinally, we describe the loss functions used for training in Section 3.5.
@@ 3.1 Problem Setup
    2D edges: $\mathcal{E}=\{E_{ij}\mid E_{ij}\in\mathbb{R}^{M_{E}\times 3}\}$, where for patch $P_{i}$ and its attached 3D curves $C_{j}$, $E_{ij}$ denotes the 2D flattened boundary edge corresponding to $C_{j}$.  [^1] A single 3D seam $C_{j}$ can induce multiple 2D edges $\{E_{ij}^{k}\}$ across its incident patches. For simplicity, we omit $k$ in the following discussion.
@@
Together, these elements provide a unified multi-level representation. The 3D patches and curves support tasks such as 3D structural perception and shape analysis, while the 2D edges and patch–curve connectivity enable 2D asset extraction for garment pattern design and physics-based simulation (via 2D triangulation of panel interiors bounded by $\{E_{ij}\}$).
@@ 4.2 Experimental Setup
Evaluation Metrics. To quantitatively evaluate our sewing pattern predictions, we define a set of evaluation metrics and also adopt several indicators from previous works [35, 25]. To assess the accuracy of topological prediction, we report: 1) Panel Count Accuracy ($\text{Acc}_{p}$), the percentage of sewing patterns with a correctly predicted number of panels; 2) Edge Count Accuracy ($\text{Acc}_{e}$), the percentage of correctly predicted edge counts within each correctly predicted panel; and 3) Overall Accuracy ($\text{Acc}_{o}$), defined as the product of Panel Count Accuracy and Edge Count Accuracy, providing a comprehensive measure of garment topology reconstruction quality. To evaluate 2D geometric precision, we report: 4) 2D Edge CD ($\text{CD}_{e}$), the Chamfer Distance between the closed-loop edge point sets of the predicted and ground-truth panels; and 5) 2D Panel IoU (IoU), the pixel-wise Intersection over Union between rasterized images of the predicted and ground-truth panels. To evaluate 3D geometric accuracy, we report: 6) Patch Point Cloud CD ($\text{CD}_{p}$), the Chamfer Distance between the predicted and ground-truth 3D patch point sets; and 7) 3D Curve CD ($\text{CD}_{c}$), the Chamfer Distance between the predicted and ground-truth 3D curve point sets. These metrics are representation-agnostic and can be applied to both our method and the baselines.
@@ 4.3 Results
2D Sewing Pattern Reconstruction. We choose AIpparel [35] as our baseline model. For fairness, we replace its monocular image input with a four-view image input. Since both GCD-TS and the GCD-MM used by AIpparel for training essentially follow the same GCD distribution, we fine-tune AIpparel on our GCD-TS dataset until the loss curve plateaus rather than training it from scratch. We refer to the fine-tuned model as AIpparel-MV (AIpparel Multi-View).
@@
Refer to captionPredicted 3D Patches/Curvesflattened 2D Edgesgeometry refinementTopology Refinement
Figure 6: Topology and geometry refinement. Without topology refinement, redundant 3D curves (red) cause incorrect 2D edges. Topology refinement removes these redundancies, producing clean panel structures, while geometry refinement enforces accurate, closed-loop boundaries suitable for triangulation.
Table 1: Quantitative Comparison of 2D Panel Quality. We compare the 2D panels generated by ReWeaver and AIpparel-MV using the metrics defined in Section 4.2. ReWeaver outperforms AIpparel-MV on five out of six metrics and achieves a comparable score on $\text{Acc}_{e}$.
@@
AIpparel-MV 0.4561 0.6774 0.3090 0.0648 0.7084
ReWeaver 0.8923 0.6570 0.5863 0.0395 0.8080
Refer to captionGround TruthReWeaverAIpparel-MV
Figure 7: Comparison with AIpparel-MV. We compare ReWeaver and AIpparel-MV against ground truth. Each example shows the predicted and ground-truth 2D panels along with the resulting simulated meshes. ReWeaver yields more accurate panels and correspondingly improved simulation results.
@@ 5 Conclusion
Limitations and Future Work.
A key limitation lies in the scarcity of high-quality 3D garments with complex topology and photorealistic textures. This gap contributes to the noticeable simulation-to-real gaps observed in our experimental inputs. Future work may alleviate this by expanding the texture library using recent advances in neural rendering or by leveraging modern generative vision models to synthesize garments with richer topological diversity.
@@
Table 3: Correction of Table 1: We present the original values and the corrected values (denoted as ReWeaver-correction) in this table.
[…]
AIpparel-MV 0.4561 0.6774 0.3090 0.0648 0.7084
ReWeaver 0.8923 0.6570 0.5863 0.0395 0.8080
ReWeaver-correction 0.9210 0.7175 0.6608 0.0391 0.8221
Table 4: Correction of Table 1: We present the original values and the corrected values (denoted as ReWeaver-correction) in this table.
[…]
AIpparel-MV 0.4561 0.6774 0.3090 0.0648 0.7084
ReWeaver-correction 0.9210 0.7175 0.6608 0.0391 0.8221
Table 5: Correction of Table 2: We present the original values and the corrected values (denoted as with refine-correction) in this table.
[…]
with refine 0.0225 0.0187 0.0264 0.8923 0.6570 0.5863 0.0395 0.8080
Typo Correction
We copied incorrect results in the main text in Table 1 and Table 2. We sincerely apologize for writing the incorrect version of the results in the main text. The corrected results for # Panel Acc, # Edge Acc, and Panel IoU are better than the previously reported version, while other results show minimal differences. This does not affect our previous analysis. We re-present the original table and the corrected results in Table 4 and Table 5.
@@ Appendix D More Results
Appendix E Discussion
@@
Unlike most previous works that focus on either 3D garment geometry or 2D panel layouts in isolation, our method reconstructs both representations jointly, which naturally enables further joint optimization to refine them simultaneously. Beyond this, one could extend our framework with additional prediction heads or differentiable cloth simulation to also estimate physical material parameters and human shape/pose from multi-view images, using the coupled 2D–3D predictions as constraints. This would allow reconstruction of more complete and better aligned 3D assets from real-world imagery, and we believe this constitutes a particularly interesting direction for future work.

### v2

@@ Abstract
High-quality 3D garment reconstruction plays a crucial role in mitigating the sim-to-real gap in applications such as digital avatars, virtual try-on and robotic manipulation. However, existing garment reconstruction methods typically rely on unstructured representations, such as 3D Gaussian Splats, struggling to provide accurate reconstructions of garment topology and sewing structures. As a result, the reconstructed outputs are often unsuitable for high-fidelity physical simulation. We propose ReWeaver, a novel framework for topology-accurate 3D garment and sewing pattern reconstruction from sparse multi-view RGB images. Given as few as four input views, ReWeaver predicts seams and panels as well as their connectivities in both the 2D UV space and the 3D space. The predicted seams and panels align precisely with the multi-view images, yielding structured 2D–3D garment representations suitable for 3D perception, high-fidelity physical simulation, and robotic manipulation. To enable effective training, we construct a large-scale dataset GCD-TS, comprising multi-view RGB images, 3D garment geometries, textured human body meshes and annotated sewing patterns. The dataset contains over 100,000 synthetic samples covering a wide range of complex geometries and topologies. Extensive experiments show that ReWeaver consistently outperforms existing methods in terms of topology accuracy, geometry alignment and seam-panel consistency. Code and data will be available at the project sii-liming.github.io/ReWeaver/.
High-quality 3D garment reconstruction plays a crucial role in mitigating the sim-to-real gap in applications such as digital avatars, virtual try-on and robotic manipulation. However, existing garment reconstruction methods typically rely on unstructured representations, such as 3D Gaussian Splats, struggling to provide accurate reconstructions of garment topology and sewing structures. As a result, the reconstructed outputs are often unsuitable for high-fidelity physical simulation. We propose ReWeaver, a novel framework for topology-accurate 3D garment and sewing pattern reconstruction from sparse multi-view RGB images. Given as few as four input views, ReWeaver predicts seams and panels as well as their connectivities in both the 2D UV space and the 3D space. The predicted seams and panels align precisely with the multi-view images, yielding structured 2D-3D garment representations suitable for 3D perception, high-fidelity physical simulation, and robotic manipulation.To enable effective training, we construct a large-scale dataset GCD-TS, comprising multi-view RGB images, 3D garment geometries, textured human body meshes and annotated sewing patterns. The dataset contains over 100,000 synthetic samples covering a wide range of complex geometries and topologies. Extensive experiments show that ReWeaver consistently outperforms existing methods in terms of topology accuracy, geometry alignment and seam-panel consistency. Code and data will be available at the project sii-liming.github.io/ReWeaver.
@@ 2.2 Optimization-based garment reconstruction
Some optimization-based methods, as well as hybrid approaches that combine learning-based prediction with physics-guided optimization, have made notable progress in recovering simulation-ready garments by integrating differentiable physics into the reconstruction process. For example, DiffAvatar [27] leverages the XPBD [33] algorithm to drive a differentiable physics engine that simulates cloth behavior, allowing gradient-based updates to align the simulated garment with the reconstructed scan geometry. However, this method relies on a predefined garment library, which limits its ability to generalize to diverse or unseen pattern structures. Furthermore, optimizing the pattern parameters requires running the differentiable simulation multiple times, making the overall process highly time-consuming. Dress-1-to-3 [26] addresses a more challenging setting by leveraging generative models as priors and using the C-IPC [23] simulation algorithm to reconstruct simulation-ready garments from a single input image. Its initial garment patterns are estimated from a single image using SewFormer [30], but these patterns have low geometric accuracy and may even contain incorrect topology. Gaussian Garments [40] reconstructs an initial static geometry from a single template frame via multi-view stereo, surface reconstruction, and remeshing, and subsequently registers it to the video sequence. However, the acquisition of the static geometry relies on multiple traditional steps, suffering from error accumulation, and such reconstruction requires the selected template frame to possess high visibility and rich textures.
@@ 3 Method
We introduce ReWeaver, a framework that reconstructs accurate geometry and topology of 3D garments from multi-view images with as few as four views. In the rest of this section, we first formally define the garment reconstruction task in Section 3.1, then introduce the building blocks of ReWeaver: a multi-view visual encoder adopted from VGGT [42] (Section 3.2); a consecutive module that predicts 3D curve and patch using a bi-path transformer (Section 3.3); and an intra-surface attention module that implicitly learns a flattening process that converts the 3D curves and patches into their 2D counterparts (Section 3.4). Finally, we describe the loss functions used for training in Section 3.5.
@@ 3.1 Problem Setup
    2D edges: $\mathcal{E}=\{E_{ij}\mid E_{ij}\in\mathbb{R}^{M_{E}\times 2}\}$, where for patch $P_{i}$ and its attached 3D curves $C_{j}$, $E_{ij}$ denotes the 2D flattened boundary edge corresponding to $C_{j}$.  [^1] A single 3D seam $C_{j}$ can induce multiple 2D edges across its incident patches. Thus, we employ two indices to denote each individual edge.
@@
Here, $N_{P}$ and $N_{C}$ denote the number of patches and curves, respectively, while $M_{P}$, $M_{C}$, and $M_{E}$ represent the number of points used to model each patch, curve, and edge. Together, these elements provide a unified multi-level representation. The 3D patches and curves support tasks such as 3D structural perception and shape analysis, while the 2D edges and patch–curve connectivity enable 2D asset extraction for garment pattern design and physics-based simulation (via 2D triangulation of panel interiors bounded by $\{E_{ij}\}$).
@@ 4.2 Experimental Setup
Evaluation Metrics. To quantitatively evaluate our sewing pattern predictions, we define a set of evaluation metrics and also adopt several indicators from previous works [36, 25]. Before computing the 2D geometric metrics, we zero-center each panel. To assess the accuracy of topological prediction, we report: 1) Panel Count Accuracy ($\text{Acc}_{p}$), the percentage of sewing patterns with a correctly predicted number of panels; 2) Edge Count Accuracy ($\text{Acc}_{e}$), the percentage of correctly predicted edge counts within each correctly predicted panel; and 3) Overall Accuracy ($\text{Acc}_{o}$), defined as the product of Panel Count Accuracy and Edge Count Accuracy, providing a comprehensive measure of garment topology reconstruction quality. To evaluate 2D geometric precision, we report: 4) 2D Edge CD ($\text{CD}_{e}$), the Chamfer Distance between the closed-loop edge point sets of the predicted and ground-truth panels.; and 5) 2D Panel IoU (IoU), the pixel-wise Intersection over Union between rasterized images of the predicted and ground-truth panels. To evaluate 3D geometric accuracy, we report: 6) Patch Point Cloud CD ($\text{CD}_{p}$), the Chamfer Distance between the predicted and ground-truth 3D patch point sets; and 7) 3D Curve CD ($\text{CD}_{c}$), the Chamfer Distance between the predicted and ground-truth 3D curve point sets. These metrics are representation-agnostic and can be applied to both our method and the baselines.
@@ 4.3 Results
2D Sewing Pattern Reconstruction. We choose Sewformer [30], ChatGarment [6], and AIpparel [36] as our baseline models. For fairness, we replace their monocular image input with a four-view image input. Since both GCD-TS and the GCD-MM used by AIpparel for training essentially follow the same GCD distribution, we fine-tune AIpparel on our GCD-TS dataset until the loss curve plateaus rather than training it from scratch. For the same reason, we fine-tune ChatGarment instead of training from scratch, but convert the JSON files of our GCD-TS dataset into the refined version following the method described in the original paper. For the training of Sewformer, we adopt the modifications introduced in  [36].
@@
Figure 6: Topology and geometry refinement. Without topology refinement, redundant 3D curves cause incorrect 2D edges (black segment in the leftmost image). Topology refinement removes these redundancies, producing clean panel structures, while geometry refinement enforces accurate, closed-loop boundaries suitable for triangulation.
Table 1: Quantitative Comparison of 2D Panel Quality. We compare the 2D panels generated by ReWeaver and baselines using the metrics defined in Section 4.2. ReWeaver outperforms on five out of six metrics.
@@
Sewformer 0.3761 0.4802 0.1806 0.1161 0.5844
ChatGarment 0.5557 0.8012 0.4452 0.0906 0.6533
AIpparel 0.4561 0.6774 0.3090 0.0648 0.7084
ReWeaver 0.9210 0.7175 0.6608 0.0391 0.8221
Refer to captionGround TruthReWeaverAIpparel
Figure 7: Comparison with AIpparel. We compare ReWeaver and AIpparel against ground truth. Each example shows the predicted and ground-truth 2D panels along with the resulting simulated meshes. ReWeaver yields more accurate panels and correspondingly improved simulation results.
@@ Appendix D More Results
Appendix E Garment Simulation
To dress the target human model, 2D garment panels are first positioned in 3D space based on their corresponding point cloud data. These panels are then imported into Marvelous Designer [34] along with the body model, where their placement is manually refined to ensure alignment with the respective body regions. Furthermore, the sewing relationships are established using the predictions from our model, which identifies and stitches together the edges corresponding to the same 3D curves. Subsequently, we employ the software’s built-in physics engine with default material parameters to simulate the draping process. This wraps the panels around the body under collision constraints, achieving a final dressed state that conforms to the target pose. The associated project files will be released in our code repository.
Appendix F Discussion
@@
Unlike most previous works that focus on either 3D garment geometry or 2D panel layouts in isolation, our method reconstructs both representations jointly, which naturally enables further joint optimization to refine them simultaneously. Beyond this, one could extend our framework with additional prediction heads or differentiable cloth simulation to also estimate physical material parameters and human shape/pose from multi-view images, using the coupled 2D–3D predictions as constraints. This would allow reconstruction of more complete and better aligned 3D assets from real-world image, and we believe this constitutes a particularly interesting direction for future work.
Appendix G Limitations
A key limitation lies in the scarcity of high-quality 3D garments with complex topology and photorealistic textures. This limits the model’s generalization ability on out-of-distribution and real-world data. Future work could introduce greater variations in camera parameters, human shapes and poses, and backgrounds to generate more diverse synthetic data, and incorporate real-world data during training to narrow the sim-to-real gap.